# An Elliott intertwining approach to classifying actions of $\mathrm{C}^{*}$-tensor categories

### 29 Jan 2024 v2

@@ 5. Two-sided Elliott intertwining
We check the conditions of the second theorem in [12]. Our underlying category is the subcategory of $\mathrm{C}^{*}_{\mathcal{C}}$ of separable $*{}^{*}$-algebras, together with extendible cocycle morphisms (see Definition 3.17). In this category, Lemma 4.12 gives us a notion of inner automorphisms in the sense of [12]. Moreover, the topology of the morphism spaces in this category, as in Lemma 4.9, is induced by a complete metric. Indeed, let $K_{n}$ be an increasing sequence of finite sets containing $1_{\mathcal{C}}$ such that $\bigcup_{n\in\mathbb{N}}K_{n}=\mathrm{Irr}(\mathcal{C})$. For any $X\in\mathrm{Irr}(\mathcal{C})$, choose a sequence of contractions $\mu_{n}^{X}$ which is dense in the unit ball of $F(X)$. Then, the assignment
@@
However, we also decide to give the technical argument in full, picking up some more general results in the process. We start by introducing the setup for performing approximate intertwining arguments in the spirit of [40, Definition 3.3].
@@
[…]
@@
The conditions listed above are in the spirit of Elliott approximate intertwining arguments, as seen for example in [38, Definition $2.3.1$] or [40, Definition $3.3$]. Note that applying $X=1_{\mathcal{C}}$ in conditions (i)-(iv) above, together with the (vi) recover the usual assumptions in the two-sided Elliott intertwining argument. However, to boost the Elliott intertwining in the equivariant setting, one needs to preserve the equivariant structure. This is why conditions (i)-(iv) above are phrased for irreducible elements in the category. These conditions resemble the assumptions needed in [40, Definition 3.3] in the case of a twisted group action, although recall that the topology on cocycle morphisms is different.
Moreover, (i)-(ii) encode that the diagrams in 5.1 approximately commute in the topology described in Lemma 4.9. Furthermore, in (iv) we are making crucial use of our definition of the functors $F_{n}$ and $G_{n}$ to assume that $F_{n}(X)$ and $G_{n}(X)$ have a countable dense subset for any $X\in\mathcal{C}$ and any $n\in\mathbb{N}$.
@@
Remark 5.4.
To deduce the formulae above, one can work backwards. If the cocycle conjugacy $\theta$ was to exist, the fact that $A$ and $B$ are inductive limits would force the diagrams
[…]
to commute for all $k\in\mathbb{N}$. This immediately gives (5.2). Considering the associated families of linear maps for the cocycle morphisms in (5.2), we would need a similar diagram involving the linear maps to commute, which gives (5.3). The formulae in (5.4) and (5.5) can be obtained in exactly the same fashion.
@@
We will show that the limits in (5.2) and (5.3) exist and that the pair $(\theta,\{s^{X}\}_{X\in\mathcal{C}})$ is a cocycle morphism. Firstly, note that the limit in (5.2) exists and $\theta:A\to B$ is a $*{}^{*}$-homomorphism by the general Elliott intertwining, as can be found for example in [38, Proposition 2.3.2].
We first show that the limit in (5.3) exists for all $X\in\mathrm{Irr}(\mathcal{C}).$ We can employ a similar argument to the one in [38, Proposition 2.3.2]. Let $X\in K_{n}$ and given $x_{n}\in F_{n}(X)$, by condition (iv) in Definition 5.1, we may assume $x_{n}$ is contained in $\bigcup_{m>n}(h_{n,m}^{X})^{-1}(\mathcal{F}_{m}^{X})$. Moreover, we may assume that $h_{n,m}^{X}(x_{n})\in\mathcal{F}_{m}^{X}$ for all $m$ greater than or equal to some $m_{0}$.
@@
[…]
@@
$\|(s_{m+1}^{X}\circ h_{m}^{X})(x_{m})-(l_{m+1}^{X}\circ s_{m}^{X})(x_{m})\|\leq\delta_{m}+\delta_{m+1},\quad\forall\quad x_{m}\in\mathcal{F}_{m}^{X}\subseteq F_{m}(X).$
@@
We claim that $(\theta,\{s^{X}\}_{X\in\mathrm{Irr}(\mathcal{C})})$ induces a cocycle morphism. For this, we check the conditions in Lemma 3.11. For any $n\in\mathbb{N}$, $X\in\mathrm{Irr}(\mathcal{C})$, $x_{n}\in F_{n}(X)$, and any $a_{n},a_{n}^{\prime}\in A_{n}$, as $(\phi_{n,\infty},h_{n,\infty})$ is a cocycle morphism, it follows that
$s^{X}(\phi_{n,\infty}(a_{n})\rhd h_{n,\infty}^{X}(x_{n})\lhd\phi_{n,\infty}(a_{n}^{\prime}))=s^{X}(h_{n,\infty}^{X}(a_{n}\rhd x_{n}\lhd a_{n}^{\prime})).$
Since the limit in (5.3) exists and $l_{k+1,\infty}^{X},s_{k}^{X},h_{n,k}^{X}$ are cocycle morphisms, a direct calculation shows that
$s^{X}(h_{n,\infty}^{X}(a_{n}\rhd x_{n}\lhd a_{n}^{\prime}))=\theta(\phi_{n,\infty}(a_{n}))\rhd s^{X}(h_{n,\infty}^{X}(x_{n}))\lhd\theta(\phi_{n,\infty}(a_{n}^{\prime})).$
Given $X,Y\in\mathrm{Irr}(\mathcal{C})$, consider a morphism $f\in\mathrm{Hom}(X,Y)$. Since $\mathrm{Hom}(X,Y)$ is 00 if $X\neq Y$ and $\mathrm{Hom}(X,X)\cong\mathbb{C}$, it follows immediately that $G(f)\circ s^{X}=s^{Y}\circ F(f)$.
@@
[…]
@@
By density it now follows that $I_{X,Y}\circ(s^{Y}\boxtimes s^{X})=s^{X\otimes Y}\circ J_{X,Y}$. Finally, it follows from (5.2) that for any $a\in A$, $s^{1_{\mathcal{C}}}(a)=\theta(a)$. Hence, as $\mathcal{C}$ is semisimple, $(\theta,\{s^{X}\}_{X\in\mathrm{Irr}(\mathcal{C})})$ induces a cocycle morphism (see Remark 3.12). This follows as the map $s^{X}$ will be given by a direct sum of linear maps corresponding to irreducible objects. Since the same holds for $h_{n,\infty}^{X},l_{k+1,\infty}^{X},s_{k}^{X},$ and $h_{n,k}^{X}$, and the limit preserves this decomposition, the formula in (5.3) holds for any $X\in\mathcal{C}$.
It follows in the same way that $(\kappa,\{r^{X}\}_{X\in\mathcal{C}})$ given by the formulae in (5.4) and (5.5) yields a well-defined cocycle morphism. Moreover, the fact that $\theta$ and $\kappa$ are mutually inverse isomorphisms follows from [38, Proposition 2.3.2].
@@
Using conditions (i) and (ii) of Definition 5.1, it follows that the right hand side evaluated at $x\in(h_{n,k}^{X})^{-1}(\mathcal{F}_{k}^{X})$ converges to $h_{n,\infty}(x)$ uniformly as $k$ tends to infinity. Now, taking the limit as $k$ goes to infinity in (5.7) and using condition (iv) of Definition 5.1, we get that $r^{X}\circ s^{X}\circ h_{n,\infty}^{X}=h_{n,\infty}^{X}$. Similarly, $s^{X}\circ r^{X}=\mathrm{id}_{G(X)}$, so $(\theta,\{s^{X}\}_{X\in\mathcal{C}}):(A,F,J)\to(B,G,I)$ and $(\kappa,\{r^{X}\}_{X\in\mathcal{C}}):(B,G,I)\to(A,F,J)$ are mutually inverse cocycle conjugacies. ∎
We now use Theorem 5.3 to show that if we assume that the diagrams in (5.1) commute up to approximate unitary equivalence, then there exist mutually inverse cocycle conjugacies as in Theorem 5.3. The proof follows in a similar fashion to [40, Theorem 3.6] and [38, Corollary 2.3.3].
@@
Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects. Let $(F_{n},J^{(n)}):\mathcal{C}\curvearrowright A_{n}$ and $(G_{n},I^{(n)}):\mathcal{C}\curvearrowright B_{n}$ be sequences of actions on separable $\mathrm{C}^{*}$-algebras. Let
@@
[…]
@@
(5.9) $\bigcup\limits_{k>n}(h_{n,k}^{X})^{-1}(\{f_{m,k}^{X}\}_{1\leq m\leq t_{k}^{X}})\subset F_{n}(X)\quad\mathrm{and}\quad\bigcup\limits_{k>n}(l_{n,k}^{X})^{-1}(\{g_{m,k}^{X}\}_{1\leq m\leq t_{k}^{X}})\subset G_{n}(X)$
@@
[…]
@@
$\begin{array}[]{cl}(\psi_{n+1,k+1},l_{n+1,k+1})\circ(\theta_{n},s_{n})\hfil\\ \approx_{u}&(\theta_{k},s_{k})\circ(\kappa_{k},r_{k})\circ(\theta_{k-1},s_{k-1})\circ\ldots\circ(\kappa_{n+1},r_{n+1})\circ(\theta_{n},s_{n})\\ \approx_{u}&(\theta_{k},s_{k})\circ(\phi_{n,k},h_{n,k})\\ \approx_{u}&(\zeta_{k},S_{k})\circ(\phi_{n,k},h_{n,k}).\end{array}$
Composing with $(\psi_{k+1,\infty},l_{k+1,\infty})$ yields that
@@
Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects. Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $\mathrm{C}^{*}$-algebras. Let
@@ 5.1. Intertwining through reparametrisation
If $A$ is a separable $C^{*}$-algebra and $\eta:\mathbb{N}\to\mathbb{N}$ is any map with $\lim\limits_{n\to\infty}\eta(n)=\infty$, then it induces an endomorphism $\eta^{*}$ on $A_{\infty}$ via $\eta^{*}((a_{n})_{n})=(a_{\eta(n)})_{n}$. Moreover, if a $*{}^{*}$-tensor category $\mathcal{C}$ acts on a unital $*{}^{*}$-algebra $A$ by the triple $(A,F,J)$ with $F(X)$ algebraically finitely generated for all $X\in\mathcal{C}$, it follows from Lemma 2.6 that there exists an induced action on $A_{\infty}$ by the triple $(A_{\infty},F_{\infty},J^{\infty})$ (see also Remark 2.7).
@@
Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions of a $*{}^{*}$-tensor category $\mathcal{C}$ on separable $\mathrm{C}^{*}$-algebras with $B$ unital and $G(X)$ be algebraically finitely generated as a right $B$-module for all $X\in\mathcal{C}$. Suppose that $(\phi,h):(A,F,J)\to(B_{\infty},G_{\infty},I^{\infty})$ is a cocycle morphism such that for any map $\eta:\mathbb{N}\to\mathbb{N}$ with $\lim\limits_{n\to\infty}\eta(n)=\infty$, the cocycle morphisms $(\phi,h)$ and $(\eta^{*},r)\circ(\phi,h)$ are approximately unitarily equivalent. For each $X\in\mathcal{C}$, let $(h_{n}^{X})_{n}$ be any lift of $h^{X}$. Then, for every finite set $K\subseteq\mathrm{Irr}(\mathcal{C})$ containing $1_{\mathcal{C}}$, finite sets $\mathcal{F}^{X}\subseteq F(X)$ for $X\in K$, $\epsilon>0$, and $m\in\mathbb{N}$, there is an integer $k\geq m$ such that for every integer $n\geq k$ there is a unitary $u\in B$ for which
@@
(5.19) $\max\limits_{x\in\mathcal{F}^{X}}\|u_{k}\rhd h_{n_{k}}^{X}(x)\lhd u_{k}-h_{k}^{X}(x)\|\geq\epsilon,$
for any $X\in K$, $x\in\mathcal{F}^{X}$, and every unitary $u_{k}\in B$. Let $\eta:\mathbb{N}\to\mathbb{N}$ be the map $\eta(k)=n_{k}$ whenever $k\geq m$ and $\eta(k)=1$ for $k<m$. As $\eta_{k}\geq k$ for $k\geq m$, it follows that $\lim\limits_{k\to\infty}\eta(k)=\infty$. Moreover, $(\phi,h)$ and $(\eta^{*},r)\circ(\phi,h)$ are approximately unitarily equivalent, so there exists a unitary $u\in B_{\infty}$ for which
@@
Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects. Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $\mathrm{C}^{*}$-algebras with $B$ unital and $G(X)$ be algebraically finitely generated as a right $B$-module for all $X\in\mathcal{C}$. Let $(\phi,h):(A,F,J)\to(B_{\infty},G_{\infty},I^{\infty})$ be a cocycle morphism. Then the following are equivalent:
@@
Suppose now that $(\phi,h)$ and $(\eta^{*},r)\circ(\phi,h)$ are approximately unitarily equivalent for any $\eta:\mathbb{N}\rightarrow\mathbb{N}$ such that $\eta(n)$ converges to infinity. Let $K_{n}\subseteq\mathrm{Irr}(\mathcal{C})$ be an increasing sequence of finite sets containing $1_{\mathcal{C}}$ such that $\bigcup_{n=1}^{\infty}K_{n}=\mathrm{Irr}(\mathcal{C})$, and $\mathcal{F}_{n}^{X}\subseteq F(X)$ be an increasing sequence of finite sets such that $\bigcup_{n=1}^{\infty}\mathcal{F}_{n}^{X}$ is dense in $F(X)$ for any $X\in\mathrm{Irr}(\mathcal{C})$. Recursively applying Lemma 5.8 to $K=K_{n}$, $\mathcal{F}^{X}=\mathcal{F}_{n}^{X}$, $\epsilon=\frac{1}{2^{n}}$, one may pick $k_{0}=1<k_{1}<k_{2}<\ldots$ and unitaries $u_{n}\in B$ for $n\in\mathbb{N}$ such that

### 08 Apr 2026 v3

@@ 6. Two-sided Elliott intertwining
We check the conditions of the second theorem in [13]. Our underlying category is the subcategory of $\mathrm{C}^{*}_{\mathcal{C}}$ of separable C-algebras, together with extendible cocycle morphisms (see Definition 3.17). In this category, Lemma 5.7 gives us a notion of inner automorphisms in the sense of [13]. Moreover, the topology of the morphism spaces in this category, as in Lemma 5.4, is induced by a complete metric. Indeed, let $K_{n}$ be an increasing sequence of finite sets containing $1_{\mathcal{C}}$ such that $\bigcup_{n\in\mathbb{N}}K_{n}=\mathrm{Irr}(\mathcal{C})$. For any $X\in\mathrm{Irr}(\mathcal{C})$, choose a sequence of contractions $\mu_{n}^{X}$ which is dense in the unit ball of $F(X)$. Then, the assignment
@@
However, we also decide to give the technical argument in full, picking up some more general results in the process. We start by introducing the setup for performing approximate intertwining arguments in the spirit of [41, Definition 3.3].
@@
(6.1) $\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${F_{n}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{s_{n}^{X}}$$\scriptstyle{h_{n}^{X}}$${F_{n+1}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\dots}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${G_{n}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{n}^{X}}$$\scriptstyle{l_{n}^{X}}$${G_{n+1}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{n+1}^{X}}$${\dots\quad.}$
@@
The conditions listed above are in the spirit of Elliott approximate intertwining arguments, as seen for example in [40, Definition $2.3.1$] or [41, Definition $3.3$]. Note that applying $X=1_{\mathcal{C}}$ in conditions (i)-(iv) above, together with the (vi) recover the usual assumptions in the two-sided Elliott intertwining argument. However, to boost the Elliott intertwining in the equivariant setting, one needs to preserve the equivariant structure. This is why conditions (i)-(iv) above are phrased for irreducible elements in the category. These conditions resemble the assumptions needed in [41, Definition 3.3] in the case of a twisted group action, although recall that the topology on cocycle morphisms is different.
Moreover, (i)-(ii) encode that the diagrams in 6.1 approximately commute in the topology described in Lemma 5.4. Furthermore, in (iv) we are making crucial use of our definition of the functors $F_{n}$ and $G_{n}$ to assume that $F_{n}(X)$ and $G_{n}(X)$ have a countable dense subset for any $X\in\mathcal{C}$ and any $n\in\mathbb{N}$.
@@
We will show that the limits in (6.2) and (6.3) exist and that the pair $(\theta,\{s^{X}\}_{X\in\mathcal{C}})$ is a cocycle morphism. Firstly, note that the limit in (6.2) exists and $\theta:A\to B$ is a -homomorphism by the general Elliott intertwining, as can be found for example in [40, Proposition 2.3.2].
We first show that the limit in (6.3) exists for all $X\in\mathrm{Irr}(\mathcal{C}).$ We can employ a similar argument to the one in [40, Proposition 2.3.2]. Let $X\in K_{n}$ and given $x_{n}\in F_{n}(X)$, by condition (iv) in Definition 6.1, we may assume $x_{n}$ is contained in $\bigcup_{m>n}(h_{n,m}^{X})^{-1}(\mathcal{F}_{m}^{X})$. Moreover, we may assume that $h_{n,m}^{X}(x_{n})\in\mathcal{F}_{m}^{X}$ for all $m$ greater than or equal to some $m_{0}$.
@@
(6.6) ${F_{n}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h_{n,m}^{X}}$${F_{m}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{s_{m}^{X}}$$\scriptstyle{h_{m}^{X}}$${F_{m+1}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{s_{m+1}^{X}}$${G_{m+1}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{m+1}^{X}}$$\scriptstyle{l_{m+1}^{X}}$${G_{m+2}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\quad l_{m+2,\infty}^{X}}$${G(X).}$
@@
$\|(s_{m+1}^{X}\circ h_{m}^{X})(x_{m})-(l_{m+1}^{X}\circ s_{m}^{X})(x_{m})\|\leq\delta_{m}+\delta_{m+1},\quad\forall\ x_{m}\in\mathcal{F}_{m}^{X}\subseteq F_{m}(X).$
@@
We claim that $(\theta,\{s^{X}\}_{X\in\mathrm{Irr}(\mathcal{C})})$ induces a cocycle morphism. For this, we check the conditions in Lemma 3.10. Given $X,Y\in\mathrm{Irr}(\mathcal{C})$, consider a morphism $f\in\mathrm{Hom}(X,Y)$. Since $\mathrm{Hom}(X,Y)$ is 0 if $X\neq Y$ and $\mathrm{Hom}(X,X)\cong\mathbb{C}$, it follows immediately that $G(f)\circ s^{X}=s^{Y}\circ F(f)$.
@@
$F_{n}(Y)\boxtimes F_{n}(X)$$F_{n}(X\otimes Y)$$G(Y)\boxtimes G(X)$$G(X\otimes Y)$$\scriptstyle{(l_{k+1,\infty}^{Y}\circ s_{k}^{Y}\circ h_{n,k}^{Y})\boxtimes(l_{k+1,\infty}^{X}\circ s_{k}^{X}\circ h_{n,k}^{X})}$$\scriptstyle{J_{X,Y}^{(n)}}$$\scriptstyle{l_{k+1,\infty}^{X\otimes Y}\circ s_{k}^{X\otimes Y}\circ h_{n,k}^{X\otimes Y}}$$\scriptstyle{I_{X,Y}}$
@@
By density it now follows that $I_{X,Y}\circ(s^{Y}\boxtimes s^{X})=s^{X\otimes Y}\circ J_{X,Y}$. Hence, as $\mathcal{C}$ is semisimple, $(\theta,\{s^{X}\}_{X\in\mathrm{Irr}(\mathcal{C})})$ induces a cocycle morphism (see Remark 3.12). This follows as the map $s^{X}$ will be given by a direct sum of linear maps corresponding to irreducible objects. Since the same holds for $h_{n,\infty}^{X},l_{k+1,\infty}^{X},s_{k}^{X},$ and $h_{n,k}^{X}$, and the limit preserves this decomposition, the formula in (6.3) holds for any $X\in\mathcal{C}$.
It follows in the same way that $(\kappa,\{r^{X}\}_{X\in\mathcal{C}})$ given by the formulae in (6.4) and (6.5) yields a well-defined cocycle morphism. Moreover, the fact that $\theta$ and $\kappa$ are mutually inverse isomorphisms follows from [40, Proposition 2.3.2].
@@
Using conditions (i),(ii) and (vi) of Definition 6.1, it follows that the right hand side evaluated at $x\in(h_{n,k}^{X})^{-1}(\mathcal{F}_{k}^{X})$ converges to $h_{n,\infty}(x)$ uniformly on $k$ as $k$ tends to infinity. Now, taking the limit as $k$ goes to infinity in (6.7) and using condition (iv) of Definition 6.1, we get that $r^{X}\circ s^{X}\circ h_{n,\infty}^{X}=h_{n,\infty}^{X}$. Similarly, $s^{X}\circ r^{X}=\mathrm{id}_{G(X)}$, so $(\theta,\{s^{X}\}_{X\in\mathcal{C}}):(A,F,J)\to(B,G,I)$ and $(\kappa,\{r^{X}\}_{X\in\mathcal{C}}):(B,G,I)\to(A,F,J)$ are mutually inverse cocycle conjugacies. ∎
We now use Theorem 6.3 to show that if we assume that the diagrams in (6.1) commute up to approximate unitary equivalence, then there exist mutually inverse cocycle conjugacies as in Theorem 6.3. The proof follows in a similar fashion to [41, Theorem 3.6] and [40, Corollary 2.3.3].
@@
Let $\mathcal{C}$ be a semisimple C-tensor category with countably many isomorphism classes of simple objects. Let $(F_{n},J^{(n)}):\mathcal{C}\curvearrowright A_{n}$ and $(G_{n},I^{(n)}):\mathcal{C}\curvearrowright B_{n}$ be sequences of actions on separable $\mathrm{C}^{*}$-algebras. Let
@@
(6.8) $\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${F_{n}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{s_{n}^{X}}$$\scriptstyle{h_{n}^{X}}$${F_{n+1}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\dots}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${G_{n}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{n}^{X}}$$\scriptstyle{l_{n}^{X}}$${G_{n+1}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{n+1}^{X}}$${\dots\quad.}$
@@
[…]
@@
(6.16) $\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${F_{n}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{S_{n}^{X}}$$\scriptstyle{h_{n}^{X}}$${F_{n+1}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\dots}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${G_{n}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{R_{n}^{X}}$$\scriptstyle{l_{n}^{X}}$${G_{n+1}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{R_{n+1}(X)}$${\dots\quad.}$
@@
[…]
Composing with $(\psi_{k+1,\infty},l_{k+1,\infty})$ then by 4.5 we have that
@@
Let $\mathcal{C}$ be a semisimple C-tensor category with countably many isomorphism classes of simple objects. Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $\mathrm{C}^{*}$-algebras. Let
@@ 6.1. Intertwining through reparametrisation
If AA is a separable $C^{*}$-algebra and $\eta:\mathbb{N}\to\mathbb{N}$ is any map with $\lim\limits_{n\to\infty}\eta(n)=\infty$, then it induces an endomorphism $\eta^{*}$ on $A_{\infty}$ via $\eta^{*}((a_{n})_{n})=(a_{\eta(n)})_{n}$. Moreover, if a C-tensor category $\mathcal{C}$ acts on a C-algebra AA by a finite rank triple $(A,F,J)$, it follows from Lemma 2.8 that there is an induced action on $A_{\infty}$ by the triple $(A_{\infty},F_{\infty},J^{\infty})$ (see also Remark 2.9).
@@
Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions of a C-tensor category $\mathcal{C}$ on separable $\mathrm{C}^{*}$-algebras with $B$ unital and $(G,I)$ finite rank. Suppose that $(\phi,h):(A,F,J)\to(B_{\infty},G_{\infty},I^{\infty})$ is a cocycle morphism such that for any map $\eta:\mathbb{N}\to\mathbb{N}$ with $\lim\limits_{n\to\infty}\eta(n)=\infty$, the cocycle morphisms $(\phi,h)$ and $(\eta^{*},r)\circ(\phi,h)$ are approximately unitarily equivalent. For each $X\in\mathcal{C}$, let $(h_{n}^{X})_{n}\colon F(X)\rightarrow l^{\infty}(\mathbb{N},G(X))$ be a linear lift of $h^{X}$. Then, for every finite set $K\subseteq\mathrm{Irr}(\mathcal{C})$ containing $1_{\mathcal{C}}$, finite sets $\mathcal{F}^{X}\subseteq F(X)$ for $X\in K$, $\epsilon>0$, and $m\in\mathbb{N}$, there is an integer $k\geq m$ such that for every integer $n\geq k$ there is a unitary $u\in B$ for which
@@
(6.19) $\max\limits_{x\in\mathcal{F}^{X}}\|u_{k}\rhd h_{n_{k}}^{X}(x)\lhd u_{k}^{*}-h_{k}^{X}(x)\|\geq\epsilon,$
for any $X\in K$, $x\in\mathcal{F}^{X}$, and every unitary $u_{k}\in B$. Let $\eta:\mathbb{N}\to\mathbb{N}$ be the map $\eta(k)=n_{k}$ whenever $k\geq m$ and $\eta(k)=1$ for $k<m$. As $\eta(k)\geq k$ for $k\geq m$, it follows that $\lim\limits_{k\to\infty}\eta(k)=\infty$. Moreover, $(\phi,h)$ and $(\eta^{*},r)\circ(\phi,h)$ are approximately unitarily equivalent, so there exists a unitary $u\in B_{\infty}$ for which
@@
Let $\mathcal{C}$ be a semisimple C-tensor category with countably many isomorphism classes of simple objects. Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $\mathrm{C}^{*}$-algebras with $B$ unital and $(G,I)$ finite rank. Let $(\phi,h):(A,F,J)\to(B_{\infty},G_{\infty},I^{\infty})$ be a cocycle morphism. Then the following are equivalent:
@@
Suppose now that $(\phi,h)$ and $(\eta^{*},r)\circ(\phi,h)$ are approximately unitarily equivalent for any $\eta:\mathbb{N}\rightarrow\mathbb{N}$ such that $\eta(n)$ converges to infinity. Let $K_{n}\subseteq\mathrm{Irr}(\mathcal{C})$ be an increasing sequence of finite sets containing $1_{\mathcal{C}}$ such that $\bigcup_{n=1}^{\infty}K_{n}=\mathrm{Irr}(\mathcal{C})$, and $\mathcal{F}_{n}^{X}\subseteq F(X)$ be an increasing sequence of finite sets such that $\bigcup_{n=1}^{\infty}\mathcal{F}_{n}^{X}$ is dense in $F(X)$ for any $X\in\mathrm{Irr}(\mathcal{C})$. Let $(h_{n}^{X})_{n}:F(X)\rightarrow l^{\infty}(\mathbb{N},G(X))$ be a linear lifting. Recursively applying Lemma 6.7 to $K=K_{n}$, $\mathcal{F}^{X}=\mathcal{F}_{n}^{X}$, $\epsilon=\frac{1}{2^{n}}$, one may pick $k_{0}=1<k_{1}<k_{2}<\ldots$ and unitaries $u_{n}\in B$ for $n\in\mathbb{N}$ such that